OpenStudy (anonymous):

Evaluate the following expression using the values given: Find 3x - y - 3z if x = -2, y = 1, and z = -2. -13 -1 1 13 @rimshaa

OpenStudy (anonymous):

3x-y-3z 3*-2-1-3*-2 -6-1+6 -1

OpenStudy (anonymous):

sub the values of x,y and z into the equation 3x - y - 3z =3(-2) - (1) -3(-2) = -6 -1 +6 (its +6 since a (-) and (-) multiplied is always (+)) = -1 (6-6 is 0)

OpenStudy (anonymous):

medal pls :)